OpenStudy (anonymous):

find the limit as x approaches positive two 1/x-2

OpenStudy (anonymous):

Well, as x approaches 2, the bottom approaches 0 so it gets really really small. Then 1 over that must be approaching what?

OpenStudy (anonymous):

one?

OpenStudy (anonymous):

Nope. Look at it this way. 1/1 = 1 1/0.1 = 10 1/0.01 = 100 1/0.001 = 1000 1/0.0001 = 10000 so you can see that if we have 1 over something that is getting smaller and smaller, then the whole thing is getting bigger and bigger.

OpenStudy (anonymous):

so its approaching infinity

OpenStudy (anonymous):

Exactly :)

OpenStudy (anonymous):

but what if the positive two were negative....would it be approaching negative infinity

OpenStudy (anonymous):

 Positive two itself cannot be negative ;) But I believe what you mean is, "if x approached 2 from the negative side". Then the bottom would always be negative, so indeed the fraction would approach negative infinity. If however, x approached from the other side, the fraction would approach positive infinity.